Integrative​Risk, return, and CAPM Wolff Enterprises must consider one investment project using the capital asset pricing model​ (CAPM). Relevant information is presented in the following table.

Item

Rate of return

​Beta, b

​Risk-free asset

10​%

0.00

Market portfolio

14​%

1.00

Project

0.67

a.  Calculate the required rate of return for the​ project, given its level of nondiversifiable risk.

b.  Calculate the risk premium for the​ project, given its level of nondiverisifiable risk.

Integrative​Risk, return, and CAPM Wolff Enterprises must consider one investment project using the capital asset pricing model​ (CAPM). Relevant information is presented in the following table.

Item

Rate of return

​Beta, b

​Risk-free asset

10​%

0.00

Market portfolio

14​%

1.00

Project

0.67

a.  Calculate the required rate of return for the​ project, given its level of nondiversifiable risk.

b.  Calculate the risk premium for the​ project, given its level of nondiverisifiable risk.
Answer #1

a)Using CAPM Model to find the required return from project

R = rf+ B*(rm-rf)

Risk free rate (rf)= 10%

Market return (rm) = 14%

Beta(B)= 0.67

R = 10+ 0.67*(14-10) = 12.68%

We expect return of 12.68% from project if it is nondiversifiable.

b) If it would have been diversifiable then B=1

Note: Incase of any doubt, please do comment. I will get back to you. Thanks!!

R = 10+ 1*(14-10) = 14%

So is risk premium for nondiversiable risk =14%-12.68% =1.32%